# AVA-Bench: Atomic Visual Ability Benchmark for Vision Foundation Models

### v2

@@ 4 Evaluation Pipeline of AVA-Bench
To evaluate VFMs using AVA-Bench, we adopt the established LLM-based VFM evaluation protocol, employing the standard LLaVA-style two-stage training procedure to prepare the connector and LLM for VFM evaluation (details in subsection 2.2 and Figure 5). For each AVA in AVA-Bench, we fine-tune the connector and LLM while keeping the VFM frozen. Given the modest size of the training sets per AVA (typically around 6K–10K), we employ Parameter-Efficient Fine-Tuning (PEFT) mai2025lessons ; houlsby2019parameter ; tu2023visual , specifically Low-Rank Adaptation (LoRA) hu2022lora , to mitigate potential overfitting. Subsequently, the fine-tuned model is evaluated on the corresponding AVA-specific test sets.
@@ 5 Experiments
5.1 Setup
All experiment setup follow the evaluation pipeline described in section 4. More experiment details can be found in Appendix.
Metrics for AVAs. For absolute depth and counting AVAs, we utilized a normalized mean absolute error (MAE) relative to the ground-truth. This normalization ensures that errors involving greater distances or counts, which are inherently more challenging, are proportionally penalized less severely. Localization performance is evaluated using Generalized Intersection-over-Union (GIoU rezatofighi2019generalized ), color via CIEDE2000 luo2001development , and OCR through Average Normalized Levenshtein Similarity (ANLS) biten2019scene . All other AVAs employ standard accuracy metrics. Please refer to Appendix for more metric details.
@@ 5.2 Observations and Analyses
Figure 9: Impact of bounding boxes on spatial reasoning performance. With bounding boxes, all VFMs perform perfectly in spatial AVA; without them, models with weaker localization (MiDaS, SAM) perform worse.
Bounding boxes help isolate and assess specific AVAs. Bounding boxes serve as a crucial tool for disentangling AVAs, allowing us to isolate and evaluate specific capabilities rather than compounded performance on complex tasks. For example, when assessing spatial reasoning, VFMs exhibit uniformly strong and similar performance when provided with ground-truth bounding boxes, indicating a shared understanding of object relationships. However, removing the bounding box input transforms this into a composite task requiring both localization and spatial reasoning. As shown in Figure 9, this removal leads to a significant divergence in model performance. While SigLIP-2 and DINOv2 maintain high performance, Midas and SAM perform considerably worse in this composite setting. Analyzing their localization performance independently reveals that Midas and SAM possess significantly weaker localization capabilities compared to SigLIP-2 and DINOv2. This stark contrast confirms that performance degradation on composite tasks is often attributable to deficiencies in specific, untrained or underdeveloped constituent AVAs, such as localization in the case of Midas and SAM, rather than a fundamental failure in all visual understanding capacities.
Subgroup analyses reveal exceptions hidden by aggregate metrics. Aggregate metrics can sometimes obscure important nuanced insights. To gain a deeper understanding, we conduct detailed analyses by partitioning test samples based on specific criteria (e.g., object size in localization tasks) and examining whether these subgroup trends align with overall performance. We split localization testing samples based on normalized bounding box sizes (relative to image size), where 0.1 indicates an object occupies 10% of the image area. As illustrated in Figure 10, VFMs surprisingly exhibit minimal performance differences when localizing large objects (0.3–0.5). Conversely, performance disparities amplify as object size decreases, revealing significant weaknesses in MiDas and SAM for smaller objects. More subgroup analyses for other AVAs can be found in the Appendix.
@@
The role of language supervision in VFM versatility. Language-supervised VFMs, specifically SigLIP-1/2 and AIMv2, demonstrate broad competency across AVAs (Figure 7 & Figure 6). Their consistent high ranking highlights that language supervision is crucial to developing general-purpose visual reasoning abilities.
Figure 10: Detail results for localization for overall and different splits based on the ground-truth’s bounding box size normalized by image size. For example, 0.1 means the bounding box size is 10% of the image size. Higher GIoU is better.
@@ 6 Discussion
Struggles of Non-Language-Aligned VFMs. Figure 7 and Figure 6 indicate that non-language-aligned VFMs, despite exhibiting strengths in certain vision-centric or low-to-mid-level AVAs (e.g., DINOv2 excelling in orientation and SAM leadning in color), typically underperform in most AVAs. To investigate the underlying cause of these shortcomings, we conducted a preliminary analysis focusing on the visual feature representations before and after the LLM connector. Using linear probing on max-pooled visual features from DINOv2, we observed a substantial accuracy degradation from 66.3% (pre-connector) to 25.67% (post-connector) for the fine-grained recognition AVA, aligning with findings reported in previous work finer . This sharp performance drop suggests that critical visual information is often compromised during modality alignment processes. Recent studies have shown that partially unfreezing and fine-tuning the last few layers of VFMs can enhance performance chen2024sharegpt4v ; however, this approach risks eroding the generalizability that initially makes these VFMs valuable. Alternatively, agglomerative models ranzinger2024radio ; heinrich2024radio such as RADIO-2.1 demonstrate relatively robust performance on AVA-Bench, suggesting potential in combining specialized VFMs. Nevertheless, effectively aligning non-language-aligned VFMs to language modalities without sacrificing their inherent visual strengths remains a challenge, highlighting an important avenue for future research.
@@
Platonic Representation Hypothesis Holds? Recent research suggests that large-scale training may lead VFMs toward converging onto similar representational spaces–Platonic Representation Hypothesis platonic . Our findings partially support this hypothesis: for certain AVAs such as object recognition, texture, and relative depth estimation, VFMs demonstrate universally strong performance, indicative of similar underlying visual representations irrespective of training objectives. Conversely, significant performance disparities among VFMs in most other AVAs, as revealed by our evaluation, indicate limitations to the generality of this hypothesis. Thus, our study suggests that the Platonic Representation Hypothesis might only hold in certain cases and warrants further empirical scrutiny in more diverse and challenging contexts.
@@
What Are Things Going From Here? While MLLM have demonstrated remarkable versatility, they are not universally effective in all scenarios, especially in specialized domains cheng2024domain ; liang2024comprehensive . Thus, there is a growing necessity for developing specialized MLLMs kumar2024diagnostics ; li2025visual . Currently, selecting appropriate VFMs for such customized MLLMs remains largely heuristic tong2024cambrian ; sun2025multi ; li2025openvision . Our work provides actionable insights that transform this selection process from heuristic guesswork into principled engineering. By clearly identifying AVA-specific strengths and weaknesses, practitioners can now systematically choose VFMs to precisely address the particular visual demands of targeted downstream tasks. Moreover, AVA-Bench represents a critical step towards developing next-generation VFMs by providing a systematic, diagnostic, and comprehensive evaluation framework. This benchmark enables VFM developers to accurately pinpoint specific deficiencies and implement targeted improvements, fostering the creation of more robust, versatile, and well-rounded VFMs in the future.
@@ A.1 Atomic Visual Abilities (AVAs)
As mentioned in Section 3.1, AVAs are elemental visual capabilities that can be combined to address more complex visual reasoning tasks. The definitions and representative questions for each AVA can be found in Table 1. Additional qualitative illustrations are provided in Figure 2 and Figure 5.
@@
    Compositional text-to-image (T2I) benchmarks. Studies on controllable generation motivate core visual primitives—number, colour, texture, object identity, spatial relations, and more—used to construct compositional prompts huang2023t2i ; wu2024conceptmix . These primitives form an initial pool of candidate abilities.
@@
    VQA question analysis. We employ GPT-4 to summarize the visual skills demanded by VQA questions in various commonly-used datasets (VQAv2 goyal2017making , RealWorldQA RealWorldQA , GQA ainslie2023gqa , etc.), thereby enriching the pool with abilities emphasized by real-world questions.
Intersecting these two sources yields a concise yet crucial set of AVAs. Moreover, we focus strictly on pure perceptual tasks and exclude non-perceptual reasoning skills (e.g., historical context, mathematical reasoning). We provide more related work discussion in Appendix D.
@@ A.2 Dataset Curation
Spatial Reasoning wu2025spatial ; zhang2025call . We curate 15.9K image pairs from NYU-Depth V2 silberman2012indoor (indoor scenes) and LVIS gupta2019lvis and Objects365 shao2019objects365 (open-domain scenes), all containing instance segmentation annotations. In each image, two distinct, non-overlapping objects are selected—one highlighted with a blue bounding box (reference object) and another with a red bounding box (target object). The model must identify the relative spatial position of the red box with respect to the blue box, choosing from four multiple-choice options: Left above, Left below, Right above, and Right below (Figure 11). The preprocessing steps for dataset creation are summarized below:
@@
  • Ambiguous object categories were excluded via a manually curated list of label IDs that often lack clear boundaries or meaningful depth interpretations from NYU-Depth V2 (e.g., wall, floor, ceiling, etc.). GPT-o1 was used to unify labels across datasets, identifying common object names conveying the same meaning, and merging these samples accordingly.
@@
  • For each object class and spatial position category, an 80% train and 20% test split was ensured, for uniform distribution and fair evaluation.
@@
Counting yao2025countllm ; li2024llmcount . We curate a total of 11.3K images from five datasets— VQAv2 goyal2017making , FSC-147 ranjan2021learning , CARPK hsieh2017drone , LVIS gupta2019lvis , and CrowdHuman li2022video —to evaluate object counting abilities across diverse domains, including open-domain scenes, natural objects, structured environments, and densely crowded contexts. Each image is paired with a question prompting the model to count the number of instances of a specified object category, and the model must return an integer-valued answer(Figure 12). The preprocessing steps for dataset creation are summarized below:
@@
Localization Objects365 shao2019objects365 Open 27.8K 6.9K
LVIS gupta2019lvis Open
iNaturalist-2021 van2021benchmarking Bird, Animal
DIOR li2020object Remote-Sensing
Counting VQAv2 goyal2017making Open 9K 2.4K
FSC ranjan2021learning Open
CARPK hsieh2017drone Car
Crowd Surveillance li2022video People
LVIS gupta2019lvis Open
Fine-grained CUB-200-2011 wah2011caltech Bird 7.2K 1.8K
iNaturalist-2021 van2021benchmarking Fungi,Plant, Animal
FGVC-Aircraft maji2013fine Object
Absolute Depth NYU-Depth V2 silberman2012indoor Indoor Scene 6.9K 1.8K
KITTI geiger2013vision Outdoor Scene
Relative Depth NYU-Depth V2 silberman2012indoor Indoor Scene 9.2K 2.4K
KITTI geiger2013vision Outdoor Scene
OCR COCO-Text veit2016cocotextdatasetbenchmarktext Open 8.8K 2.2K
IIIT5K  MishraBMVC12 Open
TextVQA balanced_vqa_v2 Open
Orientation EgoOrientBench jung2024right Open 6.9K 1.6K
CURE-OR Temel2018_ICMLA Indoor
Object Recognition Objects365 shao2019objects365 Open 27.9K 6.9K
LVIS gupta2019lvis Open
iNaturalist-2021 van2021benchmarking Bird, Animal
DIOR li2020object Remote-Sensing
Action Recognition MiT monfort2019moments Open 12.1K 3K
Texture DTD cimpoi14describing Open 10.6K 2.7K
Kylberg kylberg2011kylberg Open
KTH-TIPS kth-tips_dataset Open
KTH-TIPS2 article Open
Spatial Reasoning Objects365 shao2019objects365 Open 13K 2.9K
LVIS gupta2019lvis Open
NYU-Depth V2 silberman2012indoor Indoor Scene
Scene Recognition Places434 zhou2017places Open 11.1K 2.8K
AID xia2017aid Remote-Sensing
Emotion RAF-DB li2019reliable Human 11.9K 5.1K
ExpW lian2020expression Human
Color Objects365 shao2019objects365 Open 11.2K 2.8K
Total - - 173.5K 45.3K
@@
Fine-grained paul2023simple ; gu2025bioclip ; zhang2025finer . We curate a total of 9K images from five fine-grained recognition domains—Bird, Animal, Fungi, Plant, and Object—to assess species-level recognition capabilities. The dataset sources include CUB-200-2011 wah2011caltech for birds, iNat21 van2021benchmarking for animals, fungi, and plants, and FGVC Aircraft maji2013fine for objects. Each sample contains an image and a question prompting the model to identify the specific species or object type(Figure 13). Construction details are as follows:
@@
Figure 13: Examples of Fine-grained recognition AVA Samples.
OCR huang2025ocr ; liu2024ocrbench . We curate 10.9K images from three OCR datasets—COCO-Text veit2016cocotextdatasetbenchmarktext , IIIT5K MishraBMVC12 , and TextVQA balanced_vqa_v2 —each containing word-level bounding box annotations. In each image, a red bounding box highlights the word to be transcribed. The model is prompted to recognize the textual content inside the box based on visual context(Figure 14). The preprocessing steps for dataset creation are summarized below:
@@
Localization wu2024visionllm ; sapkota2025object . We curate 34.8K localization-focused image–question pairs sourced from Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote-sensing). Each image contains a single object instance, and the model is prompted to identify its location by providing the bounding box coordinates(Figure 15). The preprocessing steps for dataset creation are summarized below:
@@
Recognition fu2024blink ; zhang2025revisiting ; tu2023holistic . We curate 34.8K recognition samples from four datasets spanning diverse visual domains—Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain objects), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote sensing). These samples are derived from the same images and object instances used in the localization benchmark. However, instead of asking for bounding box prediction, the recognition task requires the model to identify the object within a visually highlighted region (Figure 16).
@@
Color wang2025harnessing ; chiu2024megacoin . We curate 14K images from Objects365 shao2019objects365 and LVIS gupta2019lvis , each contributing 7K samples. This AVA focuses on assessing color perception in natural scenes. For each image, we extract a coherent color region using the following pipeline:
@@
Action szot2024grounding ; wang2025harnessing ; xie2025efficiently . We curate a total of 15K image–question pairs from the Moments in Time monfort2019moments dataset, covering a wide range of human actions and activities. Each sample is derived from a short video clip annotated with a specific action label (Figure 17). Construction details are as follows:
@@
Figure 17: Examples of Action AVA Samples.
Emotion yang2024emollm ; li2024eald . We curate 16K image-question pairs from two large-scale facial expression datasets: RAF-DB li2019reliable and ExpW lian2020expression . These datasets consist of human portraits labeled with one of seven basic emotions: happy, sad, angry, fear, surprise, neutral, and disgust. Each image is annotated with a bounding box localizing the face of interest.
@@
Scene dai2025humanvlm ; fan2024mllm . We curate 13.9K image-question pairs from two diverse datasets: Places434 zhou2017places (open-domain) and AID xia2017aid (remote sensing). Each image is paired with a multiple-choice question, where the model selects the correct scene category from a pool of 30 randomly sampled options. The full set includes 464 unique scene classes spanning a wide range of environments(Figure 18). The preprocessing steps for dataset creation are summarized below:
@@
  • A uniform distribution was maintained by extracting exactly 30 images per scene category, preventing category imbalance and ensuring consistent representation across classes.
@@
Texture eppel2025shape ; gavrikov2024vision . To assess a VFM’s ability to distinguish fine-grained visual patterns, we curate 13.2K image-question pairs from diverse surface textures using close-up images from four open-domain datasets: DTD cimpoi14describing , Kylberg kylberg2011kylberg , KTH-TIPS kth-tips_dataset and KTH-TIPS2 article . These datasets encompass a diverse array of texture types—such as striped, aluminum foil, and zigzagged—capturing subtle visual patterns that are essential for accurate texture recognition. Each image is paired with a multiple-choice question, requiring the model to select the correct texture label from a set of options(Figure 19). The preprocessing steps for dataset creation are summarized below:
  • GPT-4o was used to unify labels across datasets, identifying common texture attributes conveying similar meanings. These unified labels were manually verified to retain only one representative for synonymous terms.
@@
Orientation yin2025multimodal ; jung2024right . To evaluate viewpoint understanding, we curate 10.1K image-question pairs from two specialized datasets: CURE-OR Temel2018_ICMLA and EgoOrientBench jung2024right . These datasets provide uncluttered images of objects captured from nine distinct orientations—front, back, left, right, top, front left, front right, back left, and back right—allowing models to learn pose-specific cues without requiring bounding boxes(Figure 20). The preprocessing steps for dataset creation are summarized below:
@@
  • The question for each pair: "What is the orientation of the toy plane in the image? Choose one from below: 1. front, 2. front right, 3. right, 4. back right, 5. back, 6. back left, 7. left, 8. front left, 9. top". These nine options represent the common orientation labels provided by both datasets.
@@
Absolute Depth xia2024large ; mi2024hierarchical ; zhang2025vision . We curate 9K image-object pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes. These datasets contain aligned RGB and depth information. Each image includes an object annotated with a bounding box, and the task requires estimating the absolute depth of the object in meters. This value is then matched against discretized ground-truth bins for evaluation(Figure 21).
@@
Relative Depth xia2024large ; mi2024hierarchical ; zhang2025vision . We curate 11.6K image-object pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes, targeting the task of identifying which of two objects in an image is closer to the camera. Each image contains two distinct objects, each annotated with a bounding box. The model is asked to compare their absolute depth and choose the object that appears closer to the camera (Figure 22).

### v3

@@ 4 Evaluation Pipeline of AVA-Bench
To evaluate VFMs using AVA-Bench, we adopt the established LLM-based VFM evaluation protocol, employing the standard LLaVA-style two-stage training procedure to prepare the connector and LLM for VFM evaluation (details in subsection 2.2 and Figure 4 (a)). For each AVA in AVA-Bench, we fine-tune the connector and LLM while keeping the VFM frozen. Given the modest size of the training sets per AVA (typically around 6K–10K), we employ Parameter-Efficient Fine-Tuning (PEFT) [63, 33, 90], specifically Low-Rank Adaptation (LoRA) [35], to mitigate potential overfitting. Subsequently, the fine-tuned model is evaluated on the corresponding AVA-specific test sets.
@@ 5 Experiments
Metrics for AVAs. For absolute depth and counting AVAs, we utilized a normalized mean absolute error (MAE) relative to the ground-truth. This normalization ensures that errors involving greater distances or counts, which are inherently more challenging, are proportionally penalized less severely. Localization performance is evaluated using Generalized Intersection-over-Union (GIoU [76]), color via CIEDE2000 [58], and OCR through Average Normalized Levenshtein Similarity (ANLS) [4]. All other AVAs employ standard accuracy metrics. Please refer to Appendix for more metric details.
@@ 5.1 Observations and Analyses
Figure 8: (a) Impact of bounding boxes on spatial reasoning performance. With bounding boxes, all VFMs perform perfectly in spatial AVA; without them, models with weaker localization (MiDaS, SAM) perform worse. (b) Localization results for different splits based on ground-truth’s bounding box sizes. 0.1 means the bounding box size is 10% of the image size. Higher GIoU is better.
Bounding boxes isolate and assess specific AVAs. Bounding boxes serve as a crucial tool for disentangling AVAs, allowing us to isolate and evaluate specific AVAs rather than compounded performance on complex tasks. For example, VFMs exhibit uniformly strong and similar performance in spatial reasoning when provided with ground-truth bounding boxes, indicating a shared strong spatial ability. However, removing the bounding boxes transforms this into a composite task requiring localization and spatial reasoning. As shown in Figure 8(a), such removal leads to substantial divergence in model performance, with rankings closely mirroring their localization ability. This contrast confirms that performance degradation on composite tasks is often attributable to deficiencies in specific underdeveloped AVAs rather than a fundamental failure in all visual capacities.
Subgroup analyses reveal exceptions hidden by aggregate metrics. Aggregate metrics can sometimes obscure nuanced insights. To gain a deeper understanding, we conduct detailed analyses by partitioning test samples based on specific criteria (e.g., object size in localization tasks) and examining whether these subgroup trends align with overall performance. We split localization testing samples based on normalized bounding box sizes (relative to image size), where 0.1 indicates an object occupies 10% of the image area. As illustrated in Figure 8 (b), VFMs surprisingly exhibit minimal performance differences when localizing large objects (0.3–0.5). Conversely, performance disparities amplify as object size decreases, revealing significant weaknesses in MiDas and SAM for smaller objects. More subgroup analyses for other AVAs can be found in the Appendix.
@@
The role of language supervision in VFM. Language-supervised VFMs, specifically SigLIP-1/2 and AIMv2, demonstrate broad competency across AVAs (Figure 6 & Figure 5). Their consistent high ranking highlights that language supervision is key to developing general-purpose visual abilities.
@@ 6 Discussion
Struggles of Non-Language-Aligned VFMs. Figure 6 and Figure 5 indicate that non-language-aligned VFMs, despite exhibiting strengths in certain vision-centric or low-to-mid-level AVAs (e.g., DINOv2 in orientation and SAM in color), typically underperform in most AVAs. To investigate the underlying cause of these shortcomings, we conducted a preliminary study on visual feature representations before and after the LLM connector. Using linear probing on max-pooled visual features from DINOv2, we observed a substantial accuracy degradation from 66.3% (pre-connector) to 25.67% (post-connector) for the fine-grained recognition AVA, aligning with findings in previous work [43]. This sharp performance drop suggests that critical visual information is often compromised during modality alignment processes. Recent studies have shown that fine-tuning the last few layers of VFMs can enhance performance [8]; however, this approach risks eroding the generalizability that initially makes these VFMs valuable. Alternatively, agglomerative models [75, 32] such as RADIO-2.1 demonstrate relatively robust performance on AVA-Bench, suggesting potential in combining specialized VFMs. Nevertheless, effectively aligning non-language-aligned VFMs to language modalities without sacrificing their inherent visual strengths remains a challenge, highlighting an important avenue for future research.
@@
Platonic Representation Hypothesis Holds? Recent research suggests that large-scale training may lead VFMs toward converging onto similar representations–Platonic Representation Hypothesis [39]. Our findings partially support this hypothesis: for certain AVAs (e.g., object recognition, texture, and relative depth estimation), VFMs demonstrate universally strong performance, indicative of similar underlying visual representations irrespective of training objectives. Conversely, significant performance disparities among VFMs in other AVAs, indicate limitations to the generality of this hypothesis. Thus, our study suggests that this hypothesis might only hold in certain cases and warrants further empirical scrutiny in more diverse and challenging contexts.
@@
What Are Things Going From Here? While MLLM have demonstrated remarkable versatility, they are not universally effective in all scenarios, especially in specialized domains [11, 54]. Thus, there is a growing necessity for developing specialized MLLMs [45, 52]. Currently, selecting appropriate VFMs for such customized MLLMs remains largely heuristic [86, 81, 52]. Our work provides actionable insights that transform this selection process from heuristic guesswork into principled engineering. By clearly identifying AVA-specific strengths and weaknesses, practitioners can now systematically choose VFMs to precisely address the particular visual demands of targeted downstream tasks. Moreover, AVA-Bench represents a critical step towards developing next-generation VFMs by providing a systematic, diagnostic, and comprehensive evaluation framework. This benchmark enables VFM developers to accurately pinpoint specific deficiencies and implement targeted improvements, fostering the creation of more robust, versatile, and well-rounded VFMs in the future.
[…]
Supplementary Material
@@
Disclosure of LLM Usage. Portions of this manuscript were polished for clarity and readability using an LLM. The LLM was not used to generate research ideas, design experiments, analyze data, or draw conclusions. All scientific content, methods, and results are the authors’ original work.
We provide details omitted in the main paper.
@@ A.1 Atomic Visual Abilities (AVAs)
As mentioned in Section 3.1, AVAs are elemental visual capabilities that can be combined to address more complex visual reasoning tasks. The definitions and representative questions for each AVA can be found in Table 1. Additional qualitative illustrations are provided in Figure 2 and Figure 4 (b).
@@
    Compositional text-to-image (T2I) benchmarks. Studies on controllable generation motivate core visual primitives—number, colour, texture, object identity, spatial relations, and more—used to construct compositional prompts [37, 99]. These primitives form an initial pool of candidate abilities.
@@
    VQA question analysis. We employ GPT-4 to summarize the visual skills demanded by VQA questions in various commonly-used datasets (VQAv2 [27], RealWorldQA [100], GQA [1], etc.), thereby enriching the pool with abilities emphasized by real-world questions.
Intersecting these two sources yields a concise yet crucial set of AVAs. Moreover, we focus strictly on pure perceptual tasks and exclude non-perceptual reasoning skills (e.g., historical context, mathematical reasoning). We provide more related work discussion in Appendix E.
@@ A.2 Dataset Curation
Spatial Reasoning [95, 110]. We curate 11.5K image pairs from NYU-Depth V2 [80] (indoor scenes) and LVIS [30] and Objects365 [78] (open-domain scenes), all containing instance segmentation annotations. In each image, two distinct, non-overlapping objects are selected—one highlighted with a blue bounding box (reference object) and another with a red bounding box (target object). The model must identify the relative spatial position of the red box with respect to the blue box, choosing from four multiple-choice options: Left above, Left below, Right above, and Right below (Figure 9). The preprocessing steps for dataset creation are summarized below:
@@
  • For every question, we have a target and a reference object with a spatial position (the relative position of the target based on the reference object). Each object class was ensured to appear in multiple spatial positions, with 40 samples per spatial position category. For each target object class in each spatial position, we ensure diversity of reference by selecting samples from 8 distinct reference classes and drawing 5 rows per reference, thereby preventing overfitting to a small set of co-occurring anchors.
@@
  • For each target object class, reference object class, and spatial position category, an 80% train and 20% test split was ensured, for uniform distribution and fair evaluation.
@@
Figure 10: Examples of Counting AVA Samples.
Counting [105, 47]. We curate a total of 13.6K images from five datasets— VQAv2 [27], FSC-147 [74], CARPK [34], LVIS [30] , and CrowdHuman [49] —to evaluate object counting abilities across diverse domains, including open-domain scenes, natural objects, structured environments, and densely crowded contexts. Each image is paired with a question prompting the model to count the number of instances of a specified object category, and the model must return an integer-valued answer(Figure 10). The preprocessing steps for dataset creation are summarized below:
Figure 11: Examples of Fine-grained recognition AVA Samples.
@@
Localization Objects365 [78] Open 27.9K 6.9K
LVIS [30] Open
iNaturalist-2021 [91] Bird, Animal
DIOR [50] Remote-Sensing
Counting VQAv2 [27] Open 10.8K 2.8K
FSC [74] Open
CARPK [34] Car
Crowd Surveillance [49] People
LVIS [30] Open
Fine-grained CUB-200-2011 [93] Bird 7.2K 1.8K
iNaturalist-2021 [91] Fungi,Plant, Animal
FGVC-Aircraft [64] Object
Absolute Depth NYU-Depth V2 [80] Indoor Scene 6.8K 1.8K
KITTI [24] Outdoor Scene
Relative Depth NYU-Depth V2 [80] Indoor Scene 9.2K 2.4K
KITTI [24] Outdoor Scene
OCR COCO-Text [92] Open 8.8K 2.2K
IIIT5K  [67] Open
TextVQA [26] Open
Orientation EgoOrientBench [40] Open 6.9K 1.6K
CURE-OR [83] Indoor
Object Recognition Objects365 [78] Open 37.9K 7K
LVIS [30] Open
iNaturalist-2021 [91] Bird, Animal
DIOR [50] Remote-Sensing
Action Recognition MiT [68] Open 12K 3K
Texture DTD [15] Open 10.6K 2.7K
Kylberg [46] Open
KTH-TIPS [84] Open
KTH-TIPS2 [65] Open
Spatial Reasoning Objects365 [78] Open 9.9K 1.6K
LVIS [30] Open
NYU-Depth V2 [80] Indoor Scene
Scene Recognition Places434 [116] Open 11.1K 2.8K
AID [101] Remote-Sensing
Emotion RAF-DB [51] Human 11.9K 5.1K
ExpW [53] Human
Color Objects365 [78] Open 11.2K 2.8K
Total - - 182.2K 44.5K
@@
Fine-grained [71, 29, 114]. We curate a total of 9K images from five fine-grained recognition domains—Bird, Animal, Fungi, Plant, and Object—to assess species-level recognition capabilities. The dataset sources include CUB-200-2011 [93] for birds, iNat21 [91] for animals, fungi, and plants, and FGVC Aircraft [64] for objects. Each sample contains an image and a question prompting the model to identify the specific species or object type(Figure 11). Construction details are as follows:
@@
OCR [38, 56]. We curate 10.9K images from three OCR datasets—COCO-Text [92], IIIT5K [67], and TextVQA [26]—each containing word-level bounding box annotations. In each image, a red bounding box highlights the word to be transcribed. The model is prompted to recognize the textual content inside the box based on visual context(Figure 12). The preprocessing steps for dataset creation are summarized below:
@@
Localization [96, 77]. We curate 34.8K localization-focused image–question pairs sourced from Objects365 [78] and LVIS [30] (open-domain), iNaturalist-2021 [91] (birds and animals), and DIOR [50] (remote-sensing). Each image contains a single object instance, and the model is prompted to identify its location by providing the bounding box coordinates(Figure 13). The preprocessing steps for dataset creation are summarized below:
@@
Recognition [22, 112, 89]. We curate 44.9K recognition samples from four datasets spanning diverse visual domains—Objects365 [78] and LVIS [30] (open-domain objects), iNaturalist-2021 [91] (birds and animals), and DIOR [50] (remote sensing). These samples are derived from the same images and object instances used in the localization benchmark. However, instead of asking for bounding box prediction, the recognition task requires the model to identify the object within a visually highlighted region (Figure 14).
@@
Color [94, 12]. We curate 14K images from Objects365 [78] and LVIS [30], each contributing 7K samples. This AVA focuses on assessing color perception in natural scenes. For each image, we extract a coherent color region using the following pipeline:
@@
Action [82, 94, 103]. We curate a total of 15K image–question pairs from the Moments in Time [68] dataset, covering a wide range of human actions and activities. Each sample is derived from a short video clip annotated with a specific action label (Figure 15). Construction details are as follows:
@@
Emotion [104, 48]. We curate 17K image-question pairs from two large-scale facial expression datasets: RAF-DB [51] and ExpW [53]. These datasets consist of human portraits labeled with one of seven basic emotions: happy, sad, angry, fear, surprise, neutral, and disgust. Each image is annotated with a bounding box localizing the face of interest.
@@
Scene [16, 20]. We curate 13.9K image-question pairs from two diverse datasets: Places434 [116] (open-domain) and AID [101] (remote sensing). Each image is paired with a multiple-choice question, where the model selects the correct scene category from a pool of 30 randomly sampled options. The final set includes 463 unique scene classes spanning a wide range of environments(Figure 16). The preprocessing steps for dataset creation are summarized below:
@@
    Figure 15: Examples of Action AVA Samples.
  • A uniform distribution was maintained by extracting exactly 30 images per scene category, preventing category imbalance and ensuring consistent representation across classes. Classes with less than 30 images were discarded.
@@
Texture [18, 23]. To assess a VFM’s ability to distinguish fine-grained visual patterns, we curate 13.2K image-question pairs from diverse surface textures using close-up images from four open-domain datasets: DTD [15] , Kylberg [46] , KTH-TIPS [84] and KTH-TIPS2-b [65]. These datasets encompass a diverse array of texture types—such as striped, aluminum foil, and zigzagged—capturing subtle visual patterns that are essential for accurate texture recognition. Each image is paired with a multiple-choice question, requiring the model to select the correct texture label from a set of options(Figure 17). The preprocessing steps for dataset creation are summarized below:
@@
Orientation [106, 40]. To evaluate viewpoint understanding, we curate 8.5K image-question pairs from two specialized datasets: CURE-OR [83] and EgoOrientBench [40] . These datasets provide uncluttered images of objects captured from nine distinct orientations—front, back, left, right, top, front left, front right, back left, and back right—allowing models to learn pose-specific cues without requiring bounding boxes(Figure 18). The preprocessing steps for dataset creation are summarized below:
@@
  • The question for each pair: ”What is the orientation of the toy plane in the image? Choose one from below: 1. front, 2. front right, 3. right, 4. back right, 5. back, 6. back left, 7. left, 8. front left, 9. top”. These nine options represent the common orientation labels provided by both datasets.
@@
Absolute Depth [102, 66, 111]. We curate 9K image-object pairs from NYU-Depth V2 [80] for indoor scenes and KITTI [24] for outdoor scenes. These datasets contain aligned RGB and depth information. Each image includes an object annotated with a bounding box, and the task requires estimating the absolute depth of the object in meters. This value is then matched against discretized ground-truth bins for evaluation(Figure 19).
@@
Relative Depth [102, 66, 111]. We curate 11.6K image-object pairs from NYU-Depth V2 [80] for indoor scenes and KITTI [24] for outdoor scenes, targeting the task of identifying which of two objects in an image is closer to the camera. Each image contains two distinct objects, each annotated with a bounding box. The model is asked to compare their absolute depth and choose the object that appears closer to the camera (Figure 20).